40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10800     Entry time: Mon Dec 15 22:40:09 2014
Author: rana 
Type: Summary 
Category: PSL 
Subject: PMC restored 

 Found that the PMC gain has been set to 5.3 dB instead of 10 dB since 9 AM this morning, with no elog entry.

SadToastFace.jpg

I also re-aligned the beam into the PMC to minimize the reflection. It was almost all in pitch.

ELOG V3.1.3-